OpenStudy (anonymous):

does the infinite series (n+1)(n+6)/n! converge or diverge? which test should I use to find out?

OpenStudy (sirm3d):

ratio test is convenient to use with factorial terms

OpenStudy (anonymous):

it doesn't cancel out though: I end up with (n+2)(n+7)*(n+1)/(n+1)!(n+1)(n+6)

OpenStudy (anonymous):

(n+1)! = n!(n+1)

OpenStudy (anonymous):

but I still have n! left in the denominator

OpenStudy (anonymous):

ok so I have just a polynomial over a polynomial with the same degree. the limit would be 1. that doesn't help me in the ratio test

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

\[\frac{(n+2)(n+7) }{n!(n+1) } \frac{ n! }{ (n+6)(n+1)}\]

OpenStudy (anonymous):

I got up till there, I just can't simplify it in any meaningful way

OpenStudy (anonymous):

the limit is zero

OpenStudy (anonymous):

how come?

OpenStudy (anonymous):

~n^2 / n^3

OpenStudy (anonymous):

aka 1/n

OpenStudy (anonymous):

!!!! yes!! I think I forgot to multiply the second (n+1) and that's why I had n^2/n^2

OpenStudy (anonymous):

so it converges because the limit is less than 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you! I have one more question: I have (1-p/k)^k and I have to find for what p does it converge

OpenStudy (anonymous):

I would be taking the limit of 1/e^p right?

OpenStudy (anonymous):

( (1-p)/k )^k ?

OpenStudy (anonymous):

(1-(p/k))^k

OpenStudy (anonymous):